OpenStudy (anonymous):

Implicit Differential I need help with part B and C please suppose that sq. rt. of x + sq. rt. of y = 10 a) find dy/dx. I answered it which is y'=-(y 1/2) / (x 1/2) b) what is the slope of the tangent to this curve which passes through (36,16)? (Confirm first that the curve does in fact, pass through this point) c) find an equation for this line

OpenStudy (anonymous):

\(\huge \sqrt{36}+\sqrt{16}=10 \) CONFIRMED....

OpenStudy (anonymous):

(36, 16) is on the implicit curve.

OpenStudy (anonymous):

your derivative is correct so just plug in those x, y values into your derivative...

OpenStudy (anonymous):

Thank you~ and for part C do I just take 6 and 4 as part of the equation?

OpenStudy (anonymous):

where did 6, 4 come from? use (36, 16) as the point. your slope comes from the derivative... once you have the slope, you can write the equation of the tangent line in point-slope form.

OpenStudy (anonymous):

oooh I got it from the sq. rt. that you showed me on how to confirm it so, after I plugged in the x and y values it becomes -8/5 which is the slope right? (:

OpenStudy (anonymous):

point slope from: \(\large y-y_0=m(x-x_0) \) where m is your slope, and \(\large (x_0,y_0) \) is the point

OpenStudy (anonymous):

y-16=-8/5(x-36)

OpenStudy (anonymous):

not quite... your slope is wrong....

OpenStudy (anonymous):

slope is not -8/5

OpenStudy (anonymous):

your derivative is correct: \(\large y=-\frac{y^{1/2}}{x^1/2}=-\frac{\sqrt y}{\sqrt x}=-\frac{\sqrt{36}}{\sqrt{16}}=-\frac{3}{2} \)

OpenStudy (anonymous):

so your slope should be -3/2

OpenStudy (anonymous):

My mistake was that I plugged it into (y1/2)/(x1/2) with the 1/2 I understand now. Thank you so much